### v1

@@ III Outlook
New applications are emerging as our understanding of moiré physics progresses. In the following, we highlight a few recent applications:
@@
Optics, photonics, and optoelectronic devices: moiré excitons have been shown to play a pivotal role in developing tunable optoelectronic devices across various heterostructures and materials Behura et al. (2021); Guo et al. (2021); Brem et al. (2020); Liu et al. (2022). Moiré excitons serve as a platform for optoelectronic modulation, offering efficient control over light absorption, emission, and transmission essential for devices like optical switches, photodetectors, and modulators, ensuring high-speed operation and low energy consumption Ramos et al. (2022); Yu et al. (2017). For example, moiré excitons enhance device efficiency in WS/graphene heterostructures by promoting efficient carrier extraction and utilization with high sensitivity and signal-to-noise ratios Trovatello et al. (2022). Integration of moiré excitons into integrated photonic circuits enables the development of compact, scalable systems for telecommunications, sensing, and quantum information processing. Specifically, MoS/FePS heterostructures have potential for new integrated photonicsDuan et al. (2022), offering multifunctional devices with advanced performance and functionality Onga et al. (2020). Novel excitonic devices exploiting strong light-matter interactions and long-range effects will result in novel excitonic transistors, photodetectors, LEDs, and solar cells with enhanced efficiency and tunability.
@@
Quantum Information Processing: The long coherence times and strong interactions of moiré excitons position moiré devices as promising candidates for quantum information processingHuang et al. (2022). Moreover, moiré excitons serve as valuable sources of quantum light, capable of generating non-classical light states like single photons and entangled photon pairs Baek et al. (2020); Yu et al. (2017). Researchers aim to leverage these characteristics for quantum computing, communication, and cryptography, harnessing coherence and entanglement properties to encode and process quantum information effectively. Their utilization in semiconductor moiré superlattices, particularly in Kekulé/moiré superlattices, offers new prospects for practical quantum devicesYe et al. (2023). Additionally, moiré excitons show potential in magnetic field sensing applications, especially in antiferromagnet-semiconductor heterostructuresOnga et al. (2020). Their sensitivity to external magnetic fields enables high-resolution and sensitive detection, making them suitable for various magnetic imaging, navigation, and magnetometry applications.
@@
Other devices and applications: moiré excitons offer significant potential for enhancing the efficiency and functionality of various energy and detection devices Ciarrocchi et al. (2022); Du et al. (2023a). Their strong light-matter interaction in energy harvesting and photovoltaics allows for efficient photon absorption, leading to higher photoconversion efficiencies Duan et al. (2022). Additionally, their tunability enables optimization across different wavelengths, making them suitable for diverse photonic applications. In broadband photodetection, MoSe/FePS heterostructures demonstrate a broad spectral response, enhancing versatility for applications like imaging and spectroscopy Onga et al. (2020). These photodetectors operate self-driven due to efficient carrier generation and transport, resulting in low power consumption and enhanced portability. Moreover,moiré exciton-based photodetectors exhibit fast response times and low noise performance, crucial for high-speed data acquisition and sensitive detection of weak optical signals in noisy environments. Enhanced light-matter interaction in moiré excitons, particularly in Kekulé/moiré superlattices and 2D TMD materials, improves optical absorption and emission properties, benefiting applications like optical sensing and imaging Ye et al. (2023). Furthermore, moiré excitons facilitate efficient hot carrier transfer in WS/graphene heterostructures, enabling rapid carrier dynamics and high-speed device operation essential for photodetectors and communication systems.
These applications and recent understanding of the spatial variation of moiré require a new paradigm for understanding the effects of spatial inhomogeneities and atomic relaxation on the electronic properties of moiré devices. A tandem effort of new advanced characterization techniques and new machinery to calculate the structural variation of the moiré potential and collective excitations is needed towards new excitonic devices. For an extensive review of the applications of moiré optoelectronics we refer the reader to Ref. Du et al. (2023b) and Ref. Ciarrocchi et al. (2022).
@@ IV Acknowledgements
We thank Lennart Klebl for his valuable input. D.M.K gratefully acknowledges support from the DFG through SPP 2244 (Project No. 443274199) and by the Max Planck-New York City Center for Nonequilibrium Quantum Phenomena. E.M. has received funding from the Deutsche Forschungsgemeinschaft (DFG) via CRC 1083. SK acknowledges support provided by the National Science Foundation through the ExpandQISE award No. 2329087 and the Massachusetts Technology Collaborative through award number 22032.

### v2

@@ III Outlook
New technological applications are emerging as our understanding of moiré physics progresses. Semiconductor moiré superlattices and Kekulé/moiré superlattices offer new prospects for practical quantum devices Ye et al. (2023). For an extensive review of the applications of moiré optoelectronics, we refer the reader to Ref. Du et al. (2023b) and Ref. Ciarrocchi et al. (2022); Qi et al. (2022). In the following, we highlight a few functional implementations in an early stage of development:
@@
Optics, photonics, and optoelectronic devices: moiré excitons have been shown to play a pivotal role in developing tunable optoelectronic devices across various heterostructures and materials Behura et al. (2021); Guo et al. (2021); Brem et al. (2020); Liu et al. (2022b). Moiré excitons serve as a platform for optoelectronic modulation, offering efficient control over light absorption, emission, and transmission essential for devices like optical switches, photodetectors, and modulators, ensuring high-speed operation and low energy consumption Ramos et al. (2022); Yu et al. (2017). For example, moiré excitons enhance device efficiency in WS2/graphene heterostructures by promoting efficient carrier extraction and utilization with high sensitivity and signal-to-noise ratios Trovatello et al. (2022). Integration of moiré excitons into photonic circuits enables the development of compact, scalable systems for telecommunications, sensing, and quantum information processing. Specifically, MoS2/FePS3 heterostructures have the potential for new integrated photonics Duan et al. (2022), offering multifunctional devices with advanced performance and functionality Onga et al. (2020). Novel excitonic devices exploiting strong light-matter interactions and long-range effects will result in novel excitonic transistors, photodetectors, LEDs, and solar cells with enhanced efficiency and tunability.
@@
Quantum Information Processing: The long coherence times and strong interactions of moiré excitons position moiré devices as promising candidates for quantum information processing Huang et al. (2022). Moreover, moiré excitons serve as valuable sources of quantum light, capable of generating non-classical light states like single photons and entangled photon pairs Baek et al. (2020); Yu et al. (2017). Thus, these microscopic properties could be leveraged toward effectively processing quantum information in quantum computing, quantum communication, and quantum cryptography applications.
@@
Other devices and applications: moiré excitons offer significant potential for enhancing the efficiency and functionality of various energy, sensing and detection devices Ciarrocchi et al. (2022); Du et al. (2023a). Their strong light-matter interaction in energy harvesting and photovoltaics allows for efficient photon absorption and higher photoconversion efficiencies Duan et al. (2022). Additionally, their tunability enables optimization across different wavelengths, making them suitable for diverse photonic applications. In broadband photodetection, MoSe2/FePS3 heterostructures demonstrate a broad spectral response, enhancing versatility for applications like imaging and spectroscopy Onga et al. (2020). These photodetectors operate self-driven due to efficient carrier generation and transport, resulting in low power consumption and enhanced portability. Moreover, moiré exciton-based photodetectors exhibit fast response times and low noise performance, crucial for high-speed data acquisition and sensitive detection of weak optical signals in noisy environments. Enhanced light-matter interaction in moiré excitons, particularly in Kekulé/moiré superlattices and 2D TMD materials, improves optical absorption and emission properties, benefiting applications like optical sensing and imaging Ye et al. (2023). Furthermore, moiré excitons facilitate efficient hot carrier transfer in WS2/graphene heterostructures, enabling rapid carrier dynamics and high-speed device operation essential for photodetectors and communication systems. Additionally, moiré excitons show potential in magnetic field sensing applications, especially in antiferromagnet-semiconductor heterostructures Onga et al. (2020). Their sensitivity to external magnetic fields enables high-resolution and sensitive detection, making them suitable for various magnetic imaging, navigation, and magnetometry applications. Finally, the engineering of strong interlayer excitonic correlation in MoSe2/hBN/WSe2 heterostructures with equal electron and hole densities can lead to condensates at low-temperatures Qi et al. (2023a); Nguyen et al. (2023). This state is characterized by perfect Coulomb drag, in which a charge current in one layer is balanced by an equal but opposite current in the other that could be deployed in dissipationless excitonic transport devices.
This review highlights current efforts to concomitantly deploy theoretical and experimental efforts to characterize the structural variation of the moiré potential and collective excitations. However, it is fundamental to keep furthering our understanding of the role of spatial inhomogeneities of the moiré potential and local atomic relaxations on the electronic properties of moiré heterostructures towards the future deployment of novel technology based on moiré devices. In particular, how unavoidable defects introduced during the fabrication of moiré heterostructures Rhodes et al. (2019) affect the properties of moiré excitons remains largely unexplored. For example, recent experimental results have identified the presence of defect traps able to localize the exciton in MoSe2/WSe2 heterobilayer Fang et al. (2023). Additionally, atomic defects can enable new electron-phonon coupling pathways in WS2/graphene heterostructures Liu et al. (2022a). First principle calculations have proposed defect engineering as a promising route towards modulating the excitonic properties of moiré heterostructures Guo et al. (2021). While defect engineering is still in its infancy, it opens an alternative approach towards engineering the properties of v.d.W. moiré heterostructures Rakib et al. (2022). Above all, it is clear that we are currently approaching a very promising future in which 2D v.d.W. moiré physics Andrei et al. (2021) gives rise to real-world technological applications.
@@ IV Acknowledgements
We thank Lennart Klebl for his valuable input. D.M.K gratefully acknowledges support from the DFG through SPP 2244 (Project No. 443274199) and by the Max Planck-New York City Center for Nonequilibrium Quantum Phenomena. E.M. has received funding from the Deutsche Forschungsgemeinschaft (DFG) via CRC 1083. S.K. acknowledges support provided by the National Science Foundation through the ExpandQISE award No. 2329067 and the Massachusetts Technology Collaborative through award number 22032.